# Bifurcations and multistability in inducible three-gene toggle switch networks

### v1

@@ III.2 Two inducers
Activity in gene regulatory networks is not always restricted to control from a single effector. Our framework allows for induction by multiple effectors, whether synthetically through additional experimentally-inputted non-metabolizable inducers [34], or through naturally-occurring coordinated inducer activity. The direct involvement of small molecule effectors is particularly important, for example, in metabolic sugar-inducible operons. Two effectors coordinate expression in these operons – the nucleotide messenger cAMP (with its concentration determined by the presence of glucose), and a carbon source [35, 69, 70]. For other natural inducer pairs, such as in cell differentiation, we are still learning how these effectors actually induce cell fates [16], and this motivates us to consider how dynamics in our three-gene toggle switch evolve in the presence of two inducers. We specifically extend to a system in which two repressors $\bar{R}_{1}$ and $\bar{R}_{2}$ are controlled by inducer concentrations $c_{1}$ and $c_{2}$, such that dynamics evolve by the differential equations
@@
Figure 8: Phase diagram for evolving inducer concentrations $c_{1}$ and $c_{2}$ (regulating activity of $\bar{R}_{1}$ and $\bar{R}_{2}$, respectively). Each color-coded region corresponds to a phase defined by a different number of fixed points (seven in dark purple, five in light purple, three in red, and one in blue). Both inducers obey Eqn. 6 for the probability of activity with the same fixed parameters $m=2$, $\epsilon=4\>k_{B}T$, $K_{A}=150$ $\mu M$, and $K_{I}=5$ $\mu M$.
@@
Figure 9: The change in bifurcation thresholds in a single inducer system as $\alpha$ rises. (A) The number of fixed points as a function of inducer concentration. The purple threshold denotes the transition between distinct tristable dynamic regimes (i.e., seven or five fixed points), and the green threshold denotes the transition between tristable and bistable dynamics (five or three fixed points, respectively). The system corresponds to that analyzed in Section  III.1, with $\alpha=1$. (B) The change in the inducer concentration threshold at which the tristable (purple) transition occurs as a function of $\alpha$. The dark purple curve corresponds to a system where $\alpha$ targets regulation of the induced repressor, $\bar{R}_{1}$. The light purple curve is for a system where $\alpha$ affects regulation of a non-induced repressor, in this case $\bar{R}_{2}$. (C) The change in the inducer concentration defining the threshold separating tristable from bistable dynamics as a function of $\alpha$. The specific colors again correspond to the different targets for perturbation bias, with dark green targeting the induced repressor and light green the non-induced repressor.
@@
The bifurcation line seen at $p_{A}(c)\approx 0.725$ in Fig. 5 for one dimension now becomes a set of two linear thresholds intersecting at approximate inducer concentration $2.3\times 10^{-5}\>M$ (i.e., the approximate probability $p_{A}(c_{1})=p_{A}(c_{2})=0.7$. This intersection defines four regions with one (monostability in blue), two (bistability in red), or three stable points. Note that analytically we can show that the thresholds separating tristable from bistable dynamics for $c_{1}$ and $c_{2}$ are equal to those in Eqn. 29 (see Appendix A). In the purple region of Fig. 8, all three stable points are most dynamically accessible as both induction probabilities approach $1$. Note that when the probability of active repressor reaches $1$ for one of the repressors, the transition points along the axis of the other repressor reflect the bifurcations for the single-inducer system of Fig. 5, as expected. Increasing $K_{d}^{I}$ for one of the inducers would shift the curve so that the intersection of its asymptote with the corresponding inducer axis occurs at a higher concentration. This trend matches what we expect from the single inducer analysis.
@@
The results observed for one and two inducers extend naturally in the three inducer regime, with tristability when all inducer concentrations are sufficiently small, monostability when all are sufficiently large, and bistability elsewhere. Similarly to how bifurcations transformed into a linear curve in two-dimensional phase space when moving from one to two regulating inducers, as in Fig. 8, the three inducer case likewise extends these curves into corresponding bifurcation planes in three-dimensions.
@@ IV Effects of induction in systems perturbed away from symmetry
So far, we have assumed for simplicity that a given repressor binds different promoter sites with the same affinity (for example, $\bar{R}_{1}$ binds the promoters for $\bar{R}_{2}$ and $\bar{R}_{3}$ expression with the same affinity such that $K_{21}=K_{31}\equiv K_{1}$), and that protein expression is repressed with equal strengths $K_{1}=K_{2}=K_{3}\equiv K$. While this symmetric setting allows us to observe the evolution of complex bifurcation diagrams, the fact remains that many biological systems are naturally perturbed away from symmetry. It is therefore useful to introduce biases into the relative repression strengths of a given protein, defined such that the model can still take a dimensionless form using the same transformations that generated Eqns. 1719 and Eqns. 2022. This allows us to determine the sensitivity of the bifurcation thresholds to various perturbations.
@@
In Section IIIA, we observed three distinct dynamic regimes, with decreasing complexity (i.e., number of fixed points) as the inducer concentration rose. Fig. 9(A) plots these shifts in dynamic regimes explicitly as changes in the number of fixed points, with the purple threshold denoting the bifurcation between distinct tristable regimes, and the green threshold denoting the bifurcation between tristable and bistable dynamics. The system assumes equal interaction strengths among all repression-binding event types, i.e., $\alpha=1$.
We now demonstrate how these bifurcations shift as we move away from symmetry. Fig. 9(B) reveals a narrow window of perturbation away from symmetry for which a seven-fixed-point regime exists. Expressed differently, the dark purple curve indicates a window of only $0.946\lesssim\alpha\lesssim 1.028$ for which a system with sufficiently high concentrations of all three repressors is equally likely to stabilize to any of the three possible stable states. The dark green curve in Fig. 9(C) shows that the threshold separating tristable from bistable dynamics is little affected by small perturbations around $\alpha=1$ and by more sizeable perturbations $\alpha<1$, for which $\bar{R}_{1}$ has a weaker affinity for the $\bar{R}_{2}$ promoter than for the $\bar{R}_{3}$ promoter. Favoring affinity to $\bar{R}_{2}$ for $\alpha>1$, however, begins to affect the threshold inducer concentration noticeably. At $\alpha\gtrsim 1.42$, the system becomes bistable regardless of inducer concentration.
Combining the information in dark purple and dark green in Fig. 9(B) and (C), we conclude that for $\alpha<0.95$ and $1.028\lesssim\alpha\lesssim 1.42$ only two dynamic regimes are possible (with the seven-fixed point regime no longer viable at any inducer concentration). The perturbation threshold beyond which tristability of any kind is no longer possible occurs at $\alpha\approx 1.42$.
Figure 10: Gene expression in the three-gene toggle switch with self-activation. (A) Network of three mutually-repressing genes $g_{i}$, each producing a protein at average concentration $R_{i}$ that can either repress expression of $R_{j\neq i}$ or activate its own expression. (B) Regulatory binding and expression at the promoter site for gene 11. Note that a comparable illustration applies to genes 22 and $3$.
@@
The light purple and green curves in Fig. 9(B) and (C) indicate the change in thresholds as $\alpha$ rises and alters the regulatory bias of $\bar{R}_{2}$. Fig. 9(B) shows that the change in this threshold as $\alpha$ increases from $1$ (such that $\bar{R}_{2}$ has a stronger binding affinity for the promoter of the induced $\bar{R}_{1}$ than for $\bar{R}_{3}$) is comparable to the change observed in the system case described by Eqns. 3436. Decreasing $\alpha$ to favor stronger binding to the promoter for the non-induced $\bar{R}_{3}$, however, dampens the system’s response to bias. The system denoted in light purple thus has a larger window of perturbation away from symmetry where the seven-fixed-point tristable regime is possible, with the window now favoring regulation of the non-induced over the induced repressor’s expression.
The light green curve in Fig. 9(C), on the other hand, indicates that perturbation bias aimed to affect regulation by the non-induced repressor $\bar{R}_{2}$ in fact has no effect on the threshold between tristability and bistability until $\alpha\gtrsim 1.45$. We conclude that, unlike the first case targeting the induced repressor, perturbations targeting a non-induced repressor only influence how complex the system’s tristable dynamics can become.
@@ V The three-gene toggle switch with self-activation
The systems analyzed thus far focus exclusively on mutual repression. Systems exhibiting a bistable toggle switch, however, can also be modeled with gene products not only repressing other genes but also stimulating their own expressions through self-activation, as shown in Fig. 10(A). Fig. 10(B) highlights protein production and regulatory binding at a given promoter site (in this case for gene $g_{1}$), where the protein produced can act either as a repressor targeting other genes or as an activator of its own expression. These dual repression and self-activation capabilities are observed across a range of gene network motifs with varying complexity, including applications in the bacteriophage lambda switch [10, 12, 47], stem cell and developmental differentiation [16], and mammalian cell cycle progression [71]. We now consider the effect of self-activation on stability in the three-gene context.
@@
The following sections assume, as in our previous baseline model from Section IIA, that different repressors can bind non-exclusively at a given gene promoter site to regulate expression. There are two possible ways, however, to incorporate activator binding. First, we analyze cases in which activators compete with repressors for the promoter such that an activator cannot bind if any repressor is bound, and vice versa. Section V.1.2 then explores non-exclusive binding, where both an activator and a repressor can bind simultaneously at a given promoter site such that repressors temper the effect of activation.
@@
Figure 11: Two approaches to defining proteins as repressing or activating, depending on the function of effector binding. (A) Effector binding renders proteins inactive regardless of function because the binding site independently confers repressor vs. activator identity. (B) Effector binding directly determines whether a protein exists in a repressing or activating configuration, with a binding effect altering the protein from an active repressor to an active activator.
@@
Given that all proteins in this self-activating network can act as either repressors or activators, we are interested not only in whether they bind competitively or non-exclusively to the DNA, but also in determining what defines a protein’s activity as a repressor or an activator. To do this, we highlight two biological mechanisms, and how protein activities in each case are tuned by distinct quantitative roles for effector binding. In Fig. 11(A), it is the binding site identity alone that determines a protein’s function as a repressor or activator. For example, if a protein binds to its own gene’s promoter site then it functions as an activator, and otherwise as a repressor. The activity of activator proteins is then defined by the same effector binding event that captures repressor activity, with both protein types rendered inactive by effector binding.
@@
On the other hand, Fig. 11(B) presents a setting in which effector binding directly determines protein function. Since effector binding induces a protein conformational change (favoring expression in the inducer case), we propose in this case that effector binding alters a protein’s configuration from that of an active repressor to that of an active activator. Past precedent exists for modeling such effector-driven dual function transcription factors. For the arabinose operon, for example, the presence of arabinose induces a conformational change in AraC and thus determines its interaction with the DNA as either a repressor or an activator [72, 73, 74]. Effector molecules can also indirectly control transcription factor function by determining whether transcription factors recruit corepressors or coactivators when bound to the DNA, such as for steroid hormone receptors [75, 76].
@@
We demonstrate in the following discussion how these different interpretations of effector-regulated activity lead to distinct dynamics for competitive repressor-activator binding and in particular for non-exclusive binding. In comparing the two types of mechanisms for effector activity, we continue to define thermodynamic states and weights for a system that does not explicitly involve long-range binding through DNA looping, although this could actually occur in nature. Models incorporating such looping have been considered in simpler single operon studies through grand canonical ensemble interpretations  [54, 77], and could be adapted to our expanded structure and definition for induction. Our discussion here, however, will remain consistent with the analysis presented thus far, examining extensions of the bistable toggle switch to three genes with self-activation, and will focus specifically on the dynamics deriving specifically from our model of effector activity among repressors and activators.
@@ V.1 Competitive repressor-activator binding
Figure 12: Expression of each protein $R_{i}$ in the three-gene toggle switch with self-activation and competitive binding for repressors and activators. (A) Competitive binding with the role of effectors defined as in Fig. 11(A). (B) Competitive binding with the role of effectors defined as in Fig. 11(B).
@@
where $K^{(2)}=K_{2}/K_{1}$ and $K^{(3)}=K_{3}/K_{1}$. For simplicity, we assume equal binding affinities such that $K^{(2)}=K^{(3)}=1$. We now examine how the system’s dynamic profile evolves with increasing inducer concentration, and how the increasing influence of activator binding toward expression affects the dynamic landscape.
Figure 13: Comparison of the dynamics for the three-gene toggle switch without and with the presence of self-activation. In both settings, the Hill coefficient is $n=4$ and we set the same maximal production level. (A) Bifurcation diagrams for $\bar{R}_{1}$, $\bar{R}_{2}$, and $\bar{R}_{3}$ steady-state expressions without self-activation as a function of inducer concentration, as previously shown in Fig. 5(A). Maximal production is set at $\bar{a}=2$. (B) Fixed points at low inducer concentration. (C) Bifurcation diagrams for steady-state expression with self-activation as defined by Eqns. 4345. Rates are set at $\bar{a}=2$ and $\bar{b}=0$ for maximal production $\bar{a}+\bar{b}=2$. (D) Fixed points at low inducer concentration.
Figure 14: Bifurcation diagrams for $\bar{R}_{1}$, $\bar{R}_{2}$, and $\bar{R}_{3}$ expression at a low inducer concentration and at fixed $\bar{a}+\bar{b}=2$ as a function of increasing activating strength ($\bar{b}$). The corresponding set of fixed points are shown in three-dimensional expression space beneath the bifurcation diagrams, specifically highlighting each distinct dynamic phase represented by $(\bar{a},\bar{b})=(2,0)$ (also shown in Fig. 13(B)), $(\bar{a},\bar{b})=(0.5,1.5)$, $(\bar{a},\bar{b})=(0.48,1.52)$, and $(\bar{a},\bar{b})=(0,2)$.
V.1.1 Induction via Case (A)
We introduce self-activation in the induced setting of “Case (A)” from Fig. 12(A) to observe how it alters the dynamics of the baseline model. Given the states and weights listed, and assuming $K^{(2)}=K^{(3)}=1$ and a single inducer $c$ targeting $\bar{R}_{1}$, Eqns. 4345 become

### v2

@@ III.2 Two inducers
Activity in gene regulatory networks is not always restricted to control from a single effector. Our framework allows for induction by multiple effectors, whether synthetically through additional experimentally-inputted non-metabolizable inducers [50], or through naturally-occurring coordinated inducer activity. The direct involvement of small molecule effectors can be important, for example, in sugar-inducible genetic switches as they respond to metabolic nutrient availability  [70, 32, 28, 80, 77]. Much remains to be learned, however, about how natural allosteric effectors coordinate cell fate differentiation [38], and how existing regulators may play a role in inhibitory cross-talk, as evidenced in work on steroid hormone receptors [87]. These possibilities make our study of three-gene toggle switches in the presence of two inducers even more relevant.
We specifically extend to a system in which two repressors $\bar{R}_{1}$ and $\bar{R}_{2}$ are controlled by inducer concentrations $c_{1}$ and $c_{2}$, such that dynamics evolve by the differential equations
@@
Figure 8: Phase diagram for evolving inducer concentrations $c_{1}$ and $c_{2}$ (regulating activity of $\bar{R}_{1}$ and $\bar{R}_{2}$, respectively). Each color-coded region corresponds to a phase defined by a different number of fixed points (seven in dark purple, five in light purple, three in red, and one in blue). Both inducers follow Eqn. 6 for the probability of activity with the same fixed parameters $m=2$, $\Delta\epsilon=4\>k_{B}T$, $K_{A}=150$ $\mu M$, and $K_{I}=5$ $\mu M$.
@@
Figure 9: The change in bifurcation thresholds in a single inducer system as $\alpha$ rises. (A) The number of fixed points as a function of inducer concentration when $\alpha=1$. The purple threshold denotes the transition between distinct tristable dynamic regimes (i.e., seven or five fixed points), and the green threshold denotes the transition between tristable and bistable dynamics (five or three fixed points, respectively). The system corresponds to that analyzed in Section  III.1. (B) The change in the inducer concentration threshold at which the tristable (purple) transition occurs as a function of $\alpha$. The dark purple curve corresponds to a system where $\alpha$ targets regulation of the induced repressor, $\bar{R}_{1}$. The light purple curve is for a system where $\alpha$ affects regulation of a non-induced repressor, in this case $\bar{R}_{2}$. (C) The change in the inducer concentration defining the threshold separating tristable from bistable dynamics as a function of $\alpha$. The specific colors again correspond to the different targets for perturbation bias, with dark green targeting the induced repressor and light green the non-induced repressor.
@@
The bifurcation line seen earlier at $p_{A}(c)\approx 0.725$ in Fig. 5 now becomes a set of two linear thresholds in two dimensions, intersecting at approximate inducer concentration $2.3\times 10^{-5}\>M$ (i.e., the approximate probability $p_{A}(c_{1})=p_{A}(c_{2})=0.7$). This intersection defines four regions with one (monostability in blue), two (bistability in red), or three stable points (tristability in purple). Note that analytically we can show that the thresholds separating tristable from bistable dynamics for $c_{1}$ and $c_{2}$ are equal to those in Eqn. 22 (see Appendix B). In the purple region of Fig. 8, all three stable points become most accessible as both induction probabilities approach $1$. Note that when the probability of active repressor reaches $1$ for one of the repressors, the transition points along the axis of the other repressor reflect the bifurcations for the single-inducer system of Fig. 5, as expected. Increasing $K_{I}$ for one of the inducers would shift the curve so that the intersection of its asymptote with the corresponding inducer axis occurs at a higher concentration. This trend also matches what we expect from the single inducer analysis.
@@
The results observed for one and two inducers extend naturally in the three inducer regime, with tristability when all inducer concentrations are sufficiently small, monostability when all are sufficiently large, and bistability elsewhere. Similarly to how bifurcations transformed into a linear curve in two-dimensional phase space when moving from one to two regulating inducers, as in Fig. 8, the three inducer case likewise extends these curves into corresponding bifurcation planes in three dimensions.
@@ IV Effects of induction in systems perturbed away from symmetry
So far, we have assumed for simplicity that a given repressor binds different promoter sites with the same affinity (for example, $\bar{R}_{1}$ binds the promoters for $\bar{R}_{2}$ and $\bar{R}_{3}$ expression with the same affinity such that $K_{21}=K_{31}\equiv K_{1}$), and that protein expression is repressed with equal strengths $K_{1}=K_{2}=K_{3}\equiv K$. While this symmetry allows us to observe the evolution of complex bifurcation diagrams, many biological systems are naturally perturbed away from symmetry. We now introduce biases into the relative repression strengths of a given protein, defined such that the model can still take a dimensionless form using the transformations that generated Eqns. 1012 and Eqns. 1315. This allows us to determine the sensitivity of the bifurcation thresholds to various perturbations.
@@
In Section  III.1, we observed three distinct dynamic regimes, with decreasing complexity (i.e., number of fixed points) as the inducer concentration rose. Fig. 9(A) plots these shifts in dynamic regimes explicitly as changes in the number of fixed points, with the purple threshold denoting the bifurcation between distinct tristable regimes, and the green threshold denoting the bifurcation between tristable and bistable dynamics. The system sets equal interaction strengths among all repression-binding event types, i.e., $\alpha=1$.
We now demonstrate how these bifurcations shift as we move away from symmetry. Fig. 9(B) reveals a narrow window of perturbation away from symmetry for which a seven-fixed-point regime exists. Expressed differently, the dark purple curve indicates a window of only $0.946\lesssim\alpha\lesssim 1.028$ within which a system with sufficiently high concentrations of all three repressors is equally likely to stabilize to any of the three possible stable states. The dark green curve in Fig. 9(C) shows that the threshold separating tristable from bistable dynamics is little affected by perturbations in $\alpha\leq 1$, where $\bar{R}_{1}$ has a weaker affinity for the $\bar{R}_{2}$ promoter than for the $\bar{R}_{3}$ promoter. Favoring affinity to $\bar{R}_{2}$ for $\alpha>1$, however, begins to affect the threshold inducer concentration noticeably. At $\alpha\gtrsim 1.42$, the system becomes bistable regardless of inducer concentration.
Combining the information in dark purple and dark green in Fig. 9(B) and (C), we conclude that for $\alpha<0.95$ and $1.028\lesssim\alpha\lesssim 1.42$ only two dynamic regimes are accessible (with the seven-fixed point regime no longer viable at any inducer concentration). The perturbation threshold beyond which tristability of any kind is no longer possible occurs at $\alpha\approx 1.42$.
Figure 10: Gene expression in the three-gene toggle switch with self-activation. (A) Network of three mutually-repressing genes $g_{i}$, each producing a protein at average concentration $R_{i}$ that can either repress expression of $R_{j\neq i}$ or activate its own expression. (B) Regulatory binding and expression at the promoter site for gene 11. Comparable illustrations would apply for genes 22 and $3$.
@@
The light purple and green curves in Fig. 9(B) and (C) indicate the change in thresholds as $\alpha$ rises and alters the regulatory bias of $\bar{R}_{2}$. Fig. 9(B) shows that the change in this threshold as $\alpha$ increases from $1$, such that $\bar{R}_{2}$ has a stronger binding affinity for the promoter of the induced $\bar{R}_{1}$ than for $\bar{R}_{3}$, is comparable to the change observed in dark purple as described by Eqns. 2729. Decreasing $\alpha$ to favor stronger binding to the promoter for the non-induced $\bar{R}_{3}$, however, dampens the system’s response to bias. The system denoted in light purple thus has a larger window of perturbation away from symmetry where the seven-fixed-point tristable regime is possible, with the window now favoring regulation of the non-induced over the induced repressor’s expression.
The light green curve in Fig. 9(C), on the other hand, indicates that perturbation bias aimed to affect regulation by the non-induced repressor $\bar{R}_{2}$ in fact has no effect on the threshold between tristability and bistability until $\alpha\gtrsim 1.45$. We conclude that, unlike the first instance targeting the induced repressor, perturbations targeting a non-induced repressor generally influence only how complex the system’s tristable dynamics can become.
@@ V The three-gene toggle switch with self-activation
The systems analyzed thus far focus exclusively on mutual repression. Toggle switches, however, can also be modeled with gene products not only repressing other genes but also stimulating their own expressions through self-activation, as shown in Fig. 10(A). Fig. 10(B) highlights protein production and regulatory binding at a given promoter site (in this case for gene $g_{1}$), where the protein produced can act either as a repressor targeting other genes or as an activator of its own expression. These dual repression and self-activation capabilities are observed across a range of natural gene network motifs with varying complexity, including applications in the bacteriophage lambda switch [41, 63, 74], stem cell and developmental differentiation [38], and mammalian cell cycle progression [90]. We now consider the effect of self-activation on stability in the three-gene context.
@@
The following sections assume, as in our previous baseline model from Section II.1, that different repressors can bind non-exclusively at a given gene promoter site to regulate expression. There are two possible ways, however, to incorporate activator binding. First, we analyze cases in which activators compete with repressors for the promoter such that an activator cannot bind if any repressor is bound, and vice versa. Section V.1.2 then explores non-exclusive binding, where both an activator and a repressor can bind simultaneously at a given promoter site such that repressors moderate the effect of activation.
@@
Figure 11: Two approaches to defining proteins as repressing or activating, depending on the role of effector binding. (A) Effector binding renders proteins inactive regardless of function because the protein’s binding site to the DNA independently confers repressor vs. activator identity. (B) Effector binding directly determines whether a protein exists in a repressing or activating configuration, with binding transforming an active repressor to an active activator.
@@
Given that all proteins in this self-activating network can act as either repressors or activators, we are interested not only in whether they bind competitively or non-exclusively to the DNA, but also in determining what defines a protein’s activity as a repressor or an activator. To do this, we highlight two biological mechanisms, and how protein activities in each case are tuned by distinct quantitative roles for effector binding. In Fig. 11(A), it is the binding site identity alone that determines a protein’s function. For example, if a protein binds to its own gene’s promoter site then it functions as an activator, and otherwise as a repressor. Activator proteins are then defined by the same probabilistic effector binding that captures repressor activity, with both protein types rendered inactive by effector binding. We refer to this for the remainder of our study as Case I.
@@
On the other hand, Fig. 11(B) depicts a setting in which effector binding directly determines protein function. Since effector binding induces a protein conformational change (favoring expression in the inducer case), we propose in this case that effector binding alters a protein’s configuration from that of an active repressor to that of an active activator. We refer to this for the remainder of our study as Case II. Past precedent exists for modeling such effector-driven dual function transcription factors. For the arabinose operon, for example, the presence of arabinose induces a conformational change in AraC and thus determines its interaction with the DNA as either a repressor or an activator [18, 45, 72]. Effector molecules can also indirectly control transcription factor function by determining whether transcription factors recruit corepressors or coactivators when bound to the DNA, such as for steroid hormone receptors [76, 73].
@@
We demonstrate in the following discussion how these interpretations of effector-regulated activity lead to distinct dynamics for competitive repressor-activator binding and in particular for non-exclusive binding. In comparing the two effector mechanisms, we continue to define thermodynamic states and weights that do not explicitly involve long-range binding through DNA looping. Models incorporating such looping have been considered for simpler single operons through grand canonical ensemble interpretations  [82, 44], and could be adapted to our structure. Our discussion here, however, will continue to focus on our three gene systems, but now with self-activation.
@@ V.1 Competitive repressor-activator binding
Figure 12: Expression of each protein $R_{i}$ in the three-gene toggle switch with self-activation and competitive binding for repressors and activators. (A) Competitive binding with the role of effectors defined as in Fig. 11(A) (Case I). (B) Competitive binding with the role of effectors defined as in Fig. 11(B) (Case II).
@@
where $K^{(2)}=K_{2}/K_{1}$ and $K^{(3)}=K_{3}/K_{1}$. To simplify, we impose equal binding affinities such that $K^{(2)}=K^{(3)}=1$, and examine how the system’s dynamic landscape evolves with increasing inducer concentration and rising activating bias in expression.
Figure 13: Comparison of the dynamics for the three-gene toggle switch without and with the presence of self-activation. In both settings, the Hill coefficient is $n=4$ and we set the same maximal production level. (A) Bifurcation diagrams for $\bar{R}_{1}$, $\bar{R}_{2}$, and $\bar{R}_{3}$ steady-state expressions without self-activation as a function of inducer concentration, as previously shown in Fig. 5(A). Maximal production is set at $\bar{a}=2$. (B) Fixed points at a low inducer concentration, with labels highlighting the index-2 saddle and the induced $\bar{R}_{1}$-dominant index-1 saddles. (C) Bifurcation diagrams for steady-state expression with self-activation as defined by Eqns. 3638. Rates are set at $\bar{a}=2$ and $\bar{b}=0$ for maximal production satisfying $\bar{a}+\bar{b}=2$. (D) Fixed points at a low inducer concentration.
Figure 14: Bifurcation diagrams for $\bar{R}_{1}$, $\bar{R}_{2}$, and $\bar{R}_{3}$ expression at a low inducer concentration and at fixed $\bar{a}+\bar{b}=2$ as a function of increasing activation strength ($\bar{b}$). The corresponding sets of fixed points are shown in three-dimensional expression space beneath, specifically highlighting each distinct dynamic phase for $(\bar{a},\bar{b})=(2,0)$, $(\bar{a},\bar{b})=(0.5,1.5)$, $(\bar{a},\bar{b})=(0.48,1.52)$, and $(\bar{a},\bar{b})=(0,2)$.
V.1.1 Induction in Case I
We introduce self-activation in the induced setting of Case I from Fig. 12(A) to observe how it alters the dynamics of the baseline model. Given the states and weights listed and a single inducer $c$ targeting $\bar{R}_{1}$, Eqns. 3638 become